OpenStudy (mattduhhelper):

Anton bakes 36 cookies. He gives 3 cookies to each of 3 friends. How many cookies does Anton have left? 9 27 30 33

OpenStudy (studygurl14):

36 cookies minus 3 cookies per 3 friends 36 - 3(3) = ?

OpenStudy (mattduhhelper):

30

OpenStudy (mattduhhelper):

?

OpenStudy (sissyedgar):

3 cookies per friend when he has 3 friends means he gave away 3*3 cookies. 3*3 = 9 36 - 9 = 27

OpenStudy (mattduhhelper):

i originally thought that tbh